OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 x + 5y + 6 x + 5y – 6 –x – 5y + 6 –x – 5y – 6

ganeshie8 (ganeshie8):

take first equation, solve for z

ganeshie8 (ganeshie8):

x + 5y – z = 6

OpenStudy (anonymous):

alright//

ganeshie8 (ganeshie8):

to solve for z, u need to isolate z. so wat would u do to isolate it ?

OpenStudy (anonymous):

add z on both sides?

ganeshie8 (ganeshie8):

perfect ! do that

OpenStudy (anonymous):

alright...

ganeshie8 (ganeshie8):

wat do u get when u add z both sides

OpenStudy (anonymous):

you should explain how you did on my last question cause i understood it better xD

ganeshie8 (ganeshie8):

last q you're making baby steps... this q u are running a marathon it seems, so...

ganeshie8 (ganeshie8):

x + 5y – z = 6 +z +z

ganeshie8 (ganeshie8):

x + 5y = 6+z

ganeshie8 (ganeshie8):

next what ?

OpenStudy (anonymous):

i have no idea to be honest

ganeshie8 (ganeshie8):

ur goal is to isolate z, since 6 is attached to z, we subtract 6 both sides

ganeshie8 (ganeshie8):

x + 5y = 6+z -6 -6

ganeshie8 (ganeshie8):

x+5y-6 = z

ganeshie8 (ganeshie8):

z = x+5y-6

OpenStudy (anonymous):

ohhhhhhh okay